# Neural Network Enumerator for Atmospheric Chemical ODE

### v1

@@ Abstract
Modeling atmospheric chemistry is complex and computationally intense. Given the recent success of Deep neural networks in digital signal processing, we propose a Neural Network Enumerator for fast chemical concentration modeling. We consider atmospheric chemistry as a time-dependent Ordinary Differential Equation. To extract the hidden correlations between initial states and future time evolution, we propose ChemNNE , an Attention based Neural Network Enumerator (NNE) that can model the atmospheric chemistry as a neural ODE process. To efficiently simulate the chemical changes, we propose the sinusoidal time embedding to estimate the oscillating tendency over time. More importantly, we use the Fourier neural operator to model the ODE process for efficient computation. We also propose three physical-informed losses to supervise the training optimization. To evaluate our model, we propose a large-scale chemical dataset that can be used for neural network training and evaluation. The extensive experiments show that our approach achieves state-of-the-art performance in modeling accuracy and computational speed.
@@ 1 Introduction
In the past decades, deep learning has been proven as an efficient data-driven modeling approach that can handle large-scale datasets for complex digital signal processing, like images (Ramesh et al. (2021); Xu et al. (2023)), videos (Sauer et al. (2023)), audio (Radford et al. (2023)), and languages (Solaiman et al. (2019); OpenAI and et al. (2023); Touvron et al. (2023a, b)). It is intriguing to see if it can benefit chemical modeling problems. There are a number of approaches that use deep neural networks to resolve specific chemical problems, like Su et al. (2024); Goswami et al. (2024). In general, we can take the time evolution of the chemical compound concentration as an Ordinary Differential Equation (ODE), which can be modeled by the neural networks. The hidden chemical reactions are implicitly learned and optimized by training data. In such a way, we can build the neural networks as a surrogate model that can replace the numerical simulation for fast computation. Similar research works have been done in Wu et al. (2023); Ren et al. (2023). However, few studies have been conducted on large-scale atmospheric chemistry modeling or multiple-input-and-multiple-output chemical prediction. Therefore, in this study, we aim at an end-to-end efficient neural network that can model multiple chemical concentrations (49$\sim$300 aerosol chemical compounds) over time (one hour). Specifically, given the environmental parameters and initial chemical concentrations, not only it can predict the future states of the input chemical compounds, but it also predicts new chemicals generated in the process. We analyze the model performance by the concentration changes along the time by mean squared errors and running time. To interpret the learning ability of attention, we also analyze the key chemical components for their effects and correlations to other compounds. To the best of our knowledge, we propose the largest chemical concentration dataset that can be used for model training. Our proposed ChemNNE is also the first work using neural networks for atmospheric chemical enumerators. To summarize, our contributions are:
@@
    we propose the first neural network enumerator for atmospherical chemical modeling. It can take the initial chemical concentration and environmental parameters as input, to predict the future evoluations.
@@
    We propose to combine Fourier Neural Operator (FNO) and attention to model the signal in the spatial and frequency domain so that a fast neural ODE can be achieved for fast computation.
@@
    To better constrain the chemical process, we propose physical-informed losses, including identity loss, derivative loss and mass conservation loss to ensure that the model can predict further future steps without violating the chemistry.
@@ 2 Related Work
In atmospheric chemistry, one important aspect is to study the oxidation of volatile organic compounds. With prior knowledge of the emitted chemical compounds and the main atmospheric reaction paths, researchers analyze and discover atmospheric phenomena from a micro view, i.e., the concentrations of very low volatile organic compounds (ppt or ppb range), or a macro view, i.e., cloud generation and air pollution through the formation and growth of atmospheric particles. From the mathematical perspective, we study the problem as a time-series problem or partial differential equation (PDE) problem. Given the booming development of deep learning, we will discuss deep learning based chemical modeling in the following sections.
2.1 Deep learning for chemical modeling
The rise of deep learning in the early 2010s has significantly expanded the scope of scientific discovery processes. Benefiting from the graphic processing units and coupled with new algorithms, deep learning or artificial intelligence has proved to be helpful in integrating scientific knowledge to solve diverse tasks across different domains. AI for chemistry, one of the promising research directions, has recently attracted a lot of researchers for investigation, like molecular design, materials (Volk et al. (2023); Goodall and Lee (2020)), physics (Hayat et al. (2020); Portillo et al. (2020)), atmosphere (Bassetti et al. (2023); Boiko et al. (2023); Castruccio et al. (2014)) and climate (Lam and et al. (2023); Bodnar et al. (2024); Bi et al. (2023)), and so on. For example, Volk et al. (2023) proposes to use a reinforcement learning system for multi-step nanoparticle chemistry, which can automatically explore the core-shell semiconductor particles for a self-driven fluidic laboratory. Abramson et al. (2024) propose Alphafold3 for protein structure prediction, which uses a very large-scale Protein Data Bank (PDB) to train a diffusion-based neural network for accommodating arbitrary chemical components. Chithrananda et al. (2020) propose to utilize the pre-trained transformers to discover the molecular property, offering competitive performance on drug design and other molecular predictions. Hayat et al. (2020) proposes a conservative learning framework on multi-band galaxy photometry to learn image representation. Without or with very few labels, it can achieve comparable or even better performance than the supervised approaches. To extract information from galaxy spectra, Portillo et al. (2020) propose a variational autoencoder (VAE) to reduce the dimensionality for latent space spectra interpolation and outlier detection. Graphcast (Lam and et al. (2023)) is one of the promising weather forecasting AI models that uses a graph neural network to predict the global weather with medium-range resolution. Recently, Aurora has revolutionized many facets of climate studies by leveraging vast amounts of data and proposing a foundation model with 1.3B parameters, achieving state-of-the-art performance in weather forecasting. Different from weather prediction models, atmospheric chemical models study the micro and macro chemical dynamics with physical constraints. It can be very useful for extreme weather study, air pollution prevention and so on. Bassetti et al. Bassetti et al. (2023) propose a conditional emulator via the conditional diffusion model to understand the impact of human actions on the earth system. Fine-tuned from the large language models (OpenAI and et al. (2023)), ChemGPT (Boiko et al. (2023)) is proposed to assist in autonomous chemical design and analysis, like reaction optimization and experimental automation.
2.2 Learning time series for natural science
Given the initial chemical compound concentration and environmental parameters, we can model the overall process as a time series problem. To solve time series, there are two major categories: 1) autoregression and 2) ODE process. Scientific hypotheses often take the form of discrete objects, such as symbolic formulas in physics or chemical compounds in pharmaceuticals and atmospheres. A differentiable space can be used for efficient optimization as it is amendable to gradient-based methods. For autoregression, RNN (Hochreiter and Schmidhuber. (1997)) and attention (Vaswani et al. (2017)) are two major approaches for temporal modeling. The RNN-based methods utilize the recurrent structure and capture temporal variations implicitly via time-dependent state transitions. On the other hand, attention, or transformer-based approaches discover the global temporal dependencies to determine which time step needs to be attended more for future prediction. Specifically, Li et al. (2019); Zhou et al. (2021) are two efficient self-attention models that can achieve remarkable performance on forecasting. Wu et al. (2021) present Autoformer that computes the auto-correlations among the time series data via Fast Fourier Transform. FEDformer (Zhou et al. (2022)) further improves it by utilizing the mixture-of-expert design to ensemble the specific knowledge learned from the frequency domain. Knowing partial time series data, TimeNet proposes the TimesBlock to project the 1D data to the 2D space and discover the inter- and intra-time correlations for forecasting.
The ODE/PDE process is another efficient approach that can explicitly model the neural network as a partial differential process for optimization. The advent of neural ODEs (Chen et al. (2018)) has opened new possibilities for novel ODE solver methods in physics (Hayat et al. (2020); Portillo et al. (2020)), chemistry (Jiang et al. (2020); Su et al. (2024); Goswami et al. (2024)), biology (Lu et al. (2021); Pokkunuru et al. (2023)) and so on. There exist three primary branches of deep learning based ODE process: 1) continuous learning (Raissi et al. (2019, 2020); Geneva and Zabaras (2020); Norcliffe et al. (2020)), 2) discrete learning (Zhu et al. (2019); Bhatnagar et al. (2019); Khoo et al. (2020); Biloš et al. (2021); Massaroli et al. (2020); Yildiz et al. (2019)), and 3) neural operators (Li et al. (2020); K. et al. (2024); Li et al. (2023, 2020)). The continuous learning utilizes NNs as a domain projector that maps the data to the continuous latent space for capturing global solution patterns. The idea is to embed the physics-informed losses into the training process to ensure that the behaviors of the network follow the ODE process. The discrete learning models the nonlinear dynamics on mesh grids by leveraging convolutional or graph NNs. Neural operators treat the mapping from initial conditions to solutions at time t as an input-output supervised task, which can be learned by neural operations. For example, Fourier Neural Opeartor (Li et al. (2020)) parameterizes the integral kernel directly in Fourier space as simple multiplication, which is a significant computation reduction. Spherical Fourier Neural Opeartor (Bonev et al. (2023)) generalizes the Fourier Transform as a spherical geometry, achieving stable autoregressive on long-term forecasting. The recent advance of Implicit Neural Representations (INR) shows potential to be used for predicting the spatially continuous solution of an ODE/PDE given a spatial position query. In this work, we will study this method and combine it with existing ODE methods for further improvements.
@@ 3.1 Data preparation
The data set used to study atmospheric chemistry is modeled by ARCA box (Clusius et al. (2022)), which is a comprehensive toolbox for modeling atmospheric chemistry and aerosol processes. The model can be set up to any kinetic chemistry system, but most often the base chemistry scheme comes from the Master Chemical Mechanism (Jenkin et al. (2015, 1997); Saunders et al. (2003)). We used a subset of the MCM, augmented with the Peroxy Radical Autoxidation Mechanism (Roldin et al. (2019)), which simulates monoterpene autoxidation, crucial in producing low volatility vapors from biogenic volatile organic carbon (BVOC) emissions. The subset was selected so that it contains both biogenically and anthropogenically emitted compounds such as isoprene, monoterpenes, aromatics, alkenes and alkynes, and together with their reaction products contains altogether 3301 compounds and 9530 reactions. The kinetic chemistry system is converted to a system of ODEs using the Kinetic PreProcessor (Sandu et al. (2023)), and then incorporated in ARCA box and solved numerically to simulate the time series of the reaction products.
To create the dataset, we varied the initial concentrations of 49 precursors and three environmental conditions: temperature, relative humidity and short-wave radiation, affecting the reaction rates and photochemistry. Chemically related compounds were linked together and varied in tandem, resulting in 15 independently varied initial concentrations. To create the test data set the initial concentrations and environmental variables were either at their realistic upper and lower bounds, and additionally simulations where the environmental variables were in the middle values. This produced $8847362^{15}\times 3^{3}=$ 884736 different time series of the 3301 chemical compounds.
@@
3.2 ChemNNE for Atmospheric modeling
@@
In Eq 1, it comprises three components: an encoder $\varphi$, the linearODE $f$, and a decoder $\psi$. The dynamics of the chemical process are characterized as $z(t)$. Given initial chemical concentration $x_{0}$, we have the latent space initial state $z_{0}=\varphi(x_{0},k)$, which serves as the initial condition for the ODE. The integral operation is evaluated by the proposed linearODE, which is achieved by the Fourier Neural Operation process. The Encoder is made of INR and time-dependent attention modules for extracting hidden correlations among different chemical components, and the decoder uses the shared MLP layers and INR module again for time prediction.
3.2.1 Time embedding and Implicit Neural Representation (INR) mapping

### v3

# Neural Network Emulator for Atmospheric Chemical ODE

@@ Abstract
Modelling atmospheric chemistry is complex and computationally intense. Given the recent success of Deep neural networks in digital signal processing, we propose a Neural Network Emulator for fast chemical concentration modelling. We consider atmospheric chemistry as a time-dependent Ordinary Differential Equation. To extract the hidden correlations between initial states and future time evolution, we propose ChemNNE , an Attention based Neural Network Emulator (NNE) that can model the atmospheric chemistry as a neural ODE process. To efficiently capture temporal patterns in chemical concentration changes, we implement sinusoidal time embedding to represent periodic tendencies over time. Additionally, we leverage the Fourier neural operator to model the ODE process, enhancing computational efficiency and facilitating the learning of complex dynamical behavior. We introduce three physics-informed loss functions, targeting conservation laws and reaction rate constraints, to guide the training optimization process. To evaluate our model, we introduce a unique, large-scale chemical dataset designed for neural network training and validation, which can serve as a benchmark for future studies. The extensive experiments show that our approach achieves state-of-the-art performance in modelling accuracy and computational speed.
@@ 1 Introduction
In the past decades, deep learning has been proven as an efficient data-driven modelling approach that can handle large-scale datasets for complex digital signal processing, like images (Ramesh et al. (2021); Xu et al. (2023)), videos (Sauer et al. (2023)), audio (Radford et al. (2023)), and languages (Solaiman et al. (2019); OpenAI and et al. (2023); Touvron et al. (2023a, b)). It is intriguing to see if it can benefit chemical modelling problems. There are a number of approaches that use deep neural networks to resolve specific chemical problems, like Su et al. (2024); Goswami et al. (2024). In general, we can take the time evolution of the chemical compound concentration as an Ordinary Differential Equation (ODE), which can be modeled by neural networks. The underlying dynamics of chemical reactions are implicitly captured and optimized through training on data. In this way, we can build neural networks as surrogate models to replace numerical simulations for faster computation. The development of fast neural emulators also helps address the computational bottlenecks in large-scale, high-resolution chemical simulations. Specifically, within global climate models containing millions of grid cells, a fast neural network can accelerate the modeling process at the grid-cell level. This approach significantly reduces computation time by aggregating information from all grid cells to estimate crucial chemical processes relevant to high-resolution climate predictions, enabling near real-time, high-resolution forecasting. While similar research has been conducted in Wu et al. (2023b); Ren et al. (2023), few studies address large-scale atmospheric chemistry modelling or tackle multiple-input-and-multiple-output chemical predictions. Therefore, in this study, we aim at an end-to-end efficient neural network that can model multiple chemical concentrations (49$\sim$300 aerosol chemical compounds) over time (one hour). Specifically, given environmental parameters and initial chemical concentrations, the model can predict both the future states of these compounds and the emergence of new compounds from chemical interactions. We evaluate model performance by analyzing temporal concentration changes, using mean squared errors for accuracy and running time for computational efficiency. To interpret the learning ability of attention, we also analyze the key chemical components for their effects and correlations to other compounds. To the best of our knowledge, our proposed ChemNNE is the first neural network-based approach for atmospheric chemical emulation, addressing large-scale multi-compound predictions over time. To summarize, our contributions are:
@@
    We propose the first neural network emulator for atmospheric chemical modelling, designed to predict the future evolution of chemical concentrations based on initial chemical concentrations and environmental parameters.
@@
    We combine the Fourier Neural Operator (FNO) and attention to represent chemical signals in both spatial and frequency domains, facilitating a computationally efficient neural ODE model.
@@
    To better constrain the chemical process, we propose physics-informed losses, including identity loss, derivative loss and mass conservation loss to ensure that the model can predict further future steps without violating the chemistry.
@@ 2 Related Work
In atmospheric chemistry, a key focus is studying the oxidation of volatile organic compounds to understand atmospheric transformations. With prior knowledge of the emitted chemical compounds and the main atmospheric reaction paths, researchers analyze atmospheric phenomena from both a micro view, focusing on the concentrations of trace volatile organic compounds (in the ppt or ppb range), and a macro view, examining large-scale effects like cloud formation and air pollution driven by particle formation and growth. Together, these perspectives provide a comprehensive understanding of atmospheric processes. From the mathematical perspective, we study the problem as a time-series problem or partial differential equation (PDE) problem. Given the booming development of deep learning, we will discuss deep learning based chemical modelling in the following sections.
2.1 Deep learning for chemical modelling
The rise of deep learning in the early 2010s has significantly expanded the scope of scientific discovery processes. Advances in GPU technology and innovative algorithms have enabled deep learning and artificial intelligence to integrate scientific knowledge effectively, supporting diverse tasks across multiple domains. AI for chemistry, one of the promising research directions, has recently attracted a lot of researchers for investigation, like weather forecasting, air quality estimation, aerosol chemical interaction, and so on. Graphcast (Lam and et al. (2023)) is one of the promising weather forecasting AI models that uses a graph neural network to predict the global weather with medium-range resolution. Recently, Aurora Bodnar et al. (2024) has revolutionized many facets of climate studies by leveraging vast amounts of data and proposing a foundation model with 1.3B parameters, achieving state-of-the-art performance in weather forecasting. Unlike weather prediction models, atmospheric chemical models examine complex chemical dynamics, capturing interactions and transformations among compounds under physical constraints. It can be very useful for extreme weather study, air pollution prevention and so on. Bassetti et al. Bassetti et al. (2023) propose a conditional emulator via the conditional diffusion model to understand the impact of human actions on the earth system. In parallel, AI’s role in autonomous chemical design and optimization has gained attention. ChemGPT (Boiko et al. (2023)), fine-tuned from large language models like GPT-4 (OpenAI and et al. (2023)), facilitates reaction optimization and experimental automation. Air pollution mitigation, particularly through aerosol chemistry, is another focal area. For example, Hou et al. (2022) employs random forests to analyze PM2.5 levels during haze episodes pre- and post-COVID-19 lockdown, while Betancourt et al. (2023) uses graph machine learning to impute missing data in tropospheric ozone measurements, bridging gaps in atmospheric pollutant time series. Additional research has applied machine learning to model other pollutants, such as OH (Zhu et al. (2022)) and NO2 (Di et al. (2020)). Recently, there has been growing interest in using neural network emulators as fast surrogate models to replace computationally intensive numerical simulations in chemical modelling. Not only can they be built into chemical systems for fast simulation and analysis, but they can also be used in high-resolution global climate modelling. For instance, Bradley et al. (2022) demonstrates the effectiveness of attention models in simulating Atmospheric Chemistry Box Models, achieving speeds over two orders of magnitude faster than traditional solvers. Similarly, Irwin et al. (2022) fine-tunes transformer models to address cheminformatics tasks.  Adie et al. (2024a) utilizes autoencoders and LSTMs for global chemical model simulations, while Liu et al. (2021) introduces a 28-layer residual regression neural network for chemical transport modelling. Despite these advancements, most of these approaches are tailored toward spatiotemporal chemical modelling for short-term weather or air quality analysis. Comprehensive investigations into detailed atmospheric chemical reactions remain relatively unexplored and present a critical area for future research.
2.2 Neural autoregression and ODE for time series learning
Given the initial concentrations of chemical compounds and environmental parameters, the overall process of estimating future chemical changes can be framed as either an Ordinary Differential Equation (ODE) approach or as a general time series model, depending on the complexity of temporal dynamics involved. There are two primary approaches for solving time series problems: (1) autoregressive models and (2) ordinary differential equation (ODE)-based processes. Though only a few works are directly related to chemical ODE modelling, summarizing time series works can help establish a foundation for understanding how chemical reactions can be processed and modeled by deep learning.
For autoregression, RNN (Hochreiter and Schmidhuber. (1997)) and attention (Vaswani et al. (2017)) are two major approaches for temporal modelling. RNN-based methods capture sequential temporal variations through recurrent structures, where state transitions are inherently time-dependent. In contrast, attention-based or transformer-based approaches capture global temporal dependencies, identifying specific time steps that are most relevant for accurate future prediction. Specifically, Li et al. (2019); Zhou et al. (2021) propose efficient self-attention models that achieve strong forecasting performance by focusing on scalable attention mechanisms. Wu et al. (2021) introduces Autoformer, which leverages Fast Fourier Transform to compute auto-correlations within time series data, enhancing its ability to capture periodic patterns. FEDformer (Zhou et al. (2022)) further improves forecasting performance by incorporating a mixture-of-experts design, which ensembles specialized knowledge learned from the frequency domain to enhance the model’s ability to capture complex temporal patterns. In cases with partial time series data, TimeNet (Wu et al. (2023a)) proposes the TimesBlock, which projects 1D time series data into a 2D space to capture inter- and intra-time correlations, facilitating more nuanced temporal dependency modelling for improved forecasting.
Scientific hypotheses in fields like physics or chemistry are often expressed as discrete entities, such as symbolic formulas or chemical compounds. However, these can be represented in a differentiable space, making them suitable for efficient optimization using gradient-based methods. Different from general time series processing, neural ODE/PDE models offer explicit differential systems for optimization. The introduction of neural ODEs (Chen et al. (2018)) has unlocked new avenues for developing novel ODE solvers, with applications spanning physics (Hayat et al. (2020); Portillo et al. (2020)), chemistry (Jiang et al. (2020); Su et al. (2024); Goswami et al. (2024)), and biology (Lu et al. (2021); Pokkunuru et al. (2023)), among other fields. Deep learning-based ODE processes can be broadly categorized into three primary branches: (1) continuous learning (Raissi et al. (2019, 2020); Geneva and Zabaras (2020); Norcliffe et al. (2020)), (2) discrete learning (Zhu et al. (2019); Bhatnagar et al. (2019); Khoo et al. (2020); Biloš et al. (2021); Massaroli et al. (2020); Yildiz et al. (2019)), and (3) neural operators (Li et al. (2020); K. et al. (2024); Li et al. (2023, 2020)). In continuous learning, neural networks are used as domain projectors, mapping data into a continuous latent space to capture global solution patterns. This approach embeds physics-informed loss functions into the training process to ensure that the network’s behavior adheres to the underlying ODE dynamics. Discrete learning, on the other hand, models nonlinear dynamics on mesh grids by utilizing convolutional or graph neural networks.
2.3 Neural operator for spatial and temporal encoding
The recent progress in neural operators for differentiable signal representations (Sitzmann et al. (2020); Saragadam et al. (2022); Lindell et al. (2022)) shows promise for predicting continuous solutions to ODE/PDE systems based on spatiotemporal queries. It is known that using MLP for 1D convolution causes bias in the low frequency domain (Ramasinghe et al. (2024)). Neural operators solve it by providing frequency-aware parameterization, such that high-frequency information will not be suppressed. Siren (Sitzmann et al. (2020)) was one of the first methods to apply sinusoidal functions in neural networks for effective signal reconstruction, creating a foundation for continuous function representation. Fourier Neural Operators (FNO) (Li et al. (2020)) advance this by using the Fast Fourier Transform to map signals in the frequency domain, parameterizing the integral kernel as a straightforward multiplication in Fourier space, significantly reducing computation. It parameterizes the integral kernel directly in Fourier space as simple multiplication, which is a significant computation reduction. For time series processing, a learnable time encoding system allows for flexible alignment of temporal features, enhancing the model’s adaptability to various time scales and improving temporal resolution. For instance, Spherical Fourier Neural Opeartor (Bonev et al. (2023)) generalizes the Fourier Transform as a spherical geometry, achieving stable autoregressive on long-term forecasting. Adie et al. (2024b) and  Kurth et al. (2023) are the two most recent works that use the Fourier Neural operator for global weather forecasting and chemical modelling. In this work, we aim to explore this technique further, combining it with existing ODE methods to achieve more robust and accurate modelling.
@@ 3.1 Data preparation
To understand real-world atmospheric chemistry, we need to generate a large-scale dataset for neural network training. Environmental chambers and flow reactors are used to study the complex chemical reactions in the atmosphere under well-controlled conditions. Such experiments have significantly improved our understanding of atmospheric chemical processes and the fate of many volatile organic compounds (VOCs), including anthropogenic pollutants. This has increased the theoretical understanding of atmospheric chemistry, leading to numerical chemical schemes which can describe the time evolution of the atmospheric chemistry. To obtain the dataset, we use the ARCA-box model (Clusius et al. (2022)), a comprehensive toolbox for modelling atmospheric chemistry and aerosol processes. The model can be set up to any kinetic chemistry system, but most often the base chemistry scheme comes from the Master Chemical Mechanism (Jenkin et al. (2015, 1997); Saunders et al. (2003)). We used a subset of the MCM, augmented with the Peroxy Radical Autoxidation Mechanism (Roldin et al. (2019)), which simulates monoterpene autoxidation, crucial in producing low volatility vapors from biogenic volatile organic carbon (BVOC) emissions. The subset was selected so that it contains both biogenically and anthropogenically emitted compounds such as isoprene, monoterpenes, aromatics, alkenes and alkynes, and together with their reaction products contains altogether 3301 compounds and 9530 reactions. The kinetic chemistry system is converted to a system of ODEs using the Kinetic PreProcessor (Sandu et al. (2023)), and then incorporated in ARCA box and solved numerically to simulate the time series of the reaction products.
To create the dataset, we first selected representative initial concentrations of 49 precursor gas compounds and three environmental conditions: temperature, relative humidity and short-wave radiation, affecting the reaction rates and photochemistry. Chemically related compounds were linked together and varied in tandem, resulting in 15 independently varied groups of initial concentrations. In other words, the variation of initial concentrations of 49 input chemical compounds is determined by 15 independent factors. These factors were selected so that they produced realistic lower and upper bounds of the initial concentrations, and thus obtained $2^{15}$ combinations of the initial concentrations. Additionally, all simulations were repeated with the environmental variables in the middle values. This produced $3^{3}$ different combinations of physical factors. Hence we obtain total $8847362^{15}\times 3^{3}=$ 884736 different chemical reaction samples. For each 1-hour time series, we sample the concentrations every 5 minutes, totalling 12 time steps. During the 1-hour simulation, new chemical compounds are generated, and including the original 49 input chemical compounds, we observe a total of 3301 chemical compounds. For our experiments, we focus on 400 of the most important chemical compounds and ignore others.
@@
3.2 ChemNNE for Atmospheric modelling
@@
In Eq 1, it comprises three components: an encoder $\varphi$, the linearODE $f$, and a decoder $\psi$. The dynamics of the chemical process are characterized as $z(t)$. Given initial chemical concentration $x_{0}$, we have the latent space initial state $z_{0}=\varphi(x_{0},k)$, which serves as the initial condition for the ODE. The integral operation is evaluated by the proposed linearODE, which is achieved by the Fourier Neural Operator.
To obtain continuous, differentiable signal representation, we apply sine as a periodic activation function to the chemical input and hidden features, $x_{i}\rightarrow\phi_{i}(x_{i})=sin(W_{i}x_{i}+b),\text{where\ }W_{i}\sim\mathcal{U}(-6/\sqrt{N},6/\sqrt{N}),c\in\mathbb{R}$, as they are continuous changes in the physical world and it is memory efficient to compute the derivatives in the sine space without being constrained by discrete data samples. The initialized weight parameters ensure that the input to each sine activation is normally distributed with a standard deviation of 1. Physically, we can interpret the weights of the sinusoid function as angular frequencies while the biases are phase offsets. Applying the sine function keeps signal amplitude constant while expanding frequency bands for high-frequency modelling. The Encoder is made of sine-based MLP layers and time-dependent attention modules for extracting hidden correlations among different chemical components, and the decoder also uses the sine-based MLP layers for time prediction.
3.2.1 Sinusoid function for time embedding